# Structured Agent Distillation for Large Language Model

### v2

@@
Abstract
Large language models (LLMs) exhibit strong capabilities as decision-making agents by interleaving reasoning and actions, as seen in ReAct-style frameworks. Yet, their practical deployment is constrained by high inference costs and large model sizes. We propose Structured Agent Distillation, a framework that compresses large LLM-based general agents into smaller student models while preserving both reasoning fidelity and action consistency. Unlike standard token-level distillation, our method segments trajectories into [REASON] and [ACT] spans, applying segment-specific losses to align each component with the teacher’s behavior. This structure-aware supervision enables compact agents to better replicate the teacher’s decision process. Experiments on ALFWorld, HotPotQA-ReAct, and WebShop show that our approach consistently outperforms token-level and imitation learning baselines, achieving significant compression with minimal performance drop. Scaling and ablation results further highlight the importance of span-level alignment for efficient and deployable agents.
1 Introduction
@@
Despite their effectiveness, LLM-based general agents remain costly to deploy due to model size and inference overhead. To address this, recent work distills large agents into smaller student models. However, most approaches rely on token-level supervision hinton2015distilling ; liu2025toward ; liu2025rora ; li2025mutual ; liu2025rcr ; yang2025fairsmoe ; eldan2023tinystories ; gu2024minillm ; minixhofer2025cross ; cui2025multi ; zhang2024knowledge , which treats the agent trajectory as a flat token sequence and aligns predictions step by step—ignoring its structured composition of reasoning and action.
@@
Contributions.
  • We introduce a structured agent distillation framework that segments teacher trajectories into reasoning and action spans, and applies span-specific supervision via binary token masks—enabling fine-grained alignment beyond token-level imitation.
@@
2 Motivation
2.1 Why Token-Level Distillation Falls Short
@@
Table 1: Comparison between LLMs and LLM-based General Agents in terms of token efficiency
@@
Output Format Final text sequencee Task trajectory (reasoning steps + actions)
2.2 Motivation: Toward Structured Agent Distillation
@@
Table 2: Comparison of LLM agent training frameworks. Only our method supports all four dimensions of structured agent distillation. Tool: supports external API calls or tool use; Reasoning-Action alignment: aligns structured spans (e.g., [REASON], [ACT]); Segment-aware supervision: applies supervision across full reasoning-action sequences; Curriculum-guided training kumar2010self : uses trajectory difficulty for progressive training.
Framework Tool Reasoning-Action Segm-aware supervision Curriculum
@@
3 Structured Agent Distillation Framework
Figure 1: Structured Agent Distillation framework.
Structured Agent Distillation addresses this by segmenting teacher trajectories into reasoning (Thought) and interaction (Action/Observation) spans, each supervised independently to promote phase-specific alignment.
As shown in Figure 1, The teacher agent, given an initial observation and sampled task, produces a reasoning trajectory and corresponding action. These [REASON] and [ACT] components are extracted into a trajectory $\tau=(\texttt{reason},\texttt{action})$ used for curriculum sampling.
The student agent learns from this trajectory $\tau$ via two objectives: (1) CoT-Policy Alignment Loss $\mathcal{L}_{\text{CoT}}$ aligns the reasoning steps, and (2) Action Consistency Loss $\mathcal{L}_{\text{Act}}$ aligns the final action decision. The overall objective is $\mathcal{L}_{\text{Traj}}=\mathcal{L}_{\text{CoT}}+\mathcal{L}_{\text{Act}}$.
For a comprehensive analysis, please refer to the Appendix A.
3.1 Problem Formulation
@@
Given a teacher policy $\pi_{T}(\tau)$, the goal is to train a compact student policy $\pi_{\theta}(\tau)$ such that:
@@
$\tau^{\prime}=\texttt{[REASON]}\ r_{1}\cdots r_{k}\ \texttt{[ACT]}\ a_{1}\cdots a_{m}.$ (3)
We tokenize this as $x=\texttt{Tokenize}(\tau^{\prime})=(x_{1},x_{2},\ldots,x_{n})$, and assign each token $x_{t}$ a segment label $s_{t}\in\{\text{Reason},\text{Action}\}$, indicating its span. These labels are used to compute segment-aware losses during training.
3.2 Trajectory Segmentation
@@
$\tau=(\tau^{(r)},\tau^{(a)}),$
where $\tau^{(r)}$ denotes the reasoning span and $\tau^{(a)}$ denotes the action span. This segmentation is performed via lightweight rule-based parsing based on prompt templates consistent across tasks (see Appendix D). The segmented trajectory is then tokenized into a sequence $x=(x_{1},x_{2},\ldots,x_{T})$.
3.3 Segment-Aware Supervision via Token Masks
To enable structured supervision in Structured Agent Distillation, we construct two binary token-level masks over the tokenized teacher trajectory $x=(x_{1},x_{2},\ldots,x_{T})$:
  • Reasoning mask $m_{r}(t)\in\{0,1\}$:
  • Action mask $m_{a}(t)\in\{0,1\}$:
Each token belongs to at most one span, so we enforce:
$m_{r}(t)+m_{a}(t)\leq 1,\quad\forall t\in[1,T].$
This constraint ensures that each token is assigned to at most one functional span, avoiding supervision overlap between reasoning and action phases. These segment-aware masks allow us to decompose the distillation loss and apply span-specific supervision. Mask construction details and validation procedures are provided in Appendix D and F.
3.4 Structured Agent Distillation Objectives
To supervise student agents under Structured Agent Distillation, we decompose each teacher trajectory $\tau$ into a reasoning span $\tau^{(r)}$ and an action span $\tau^{(a)}$ (see Appendix D for segmentation rules). The full trajectory is tokenized into $x=(x_{1},\ldots,x_{T})$ and aligned with binary token masks $m_{r}(t)$ and $m_{a}(t)$, which indicate whether token $x_{t}$ belongs to the reasoning or action span, respectively. These segment-aware masks enable selective supervision over structurally distinct parts of the trajectory.
Given these segment-aware masks, the student policy $\pi_{\theta}$ is supervised through two loss terms:
(1) CoT-Policy Alignment Loss. We align the student’s predicted distribution $p_{S}(x_{t})$ with the teacher distribution $p_{T}(x_{t})$ for reasoning tokens using KL divergence:
$\mathcal{L}_{\text{CoT}}=\sum_{t=1}^{T}m_{r}(t)\cdot\text{KL}(p_{T}(x_{t})\|p_{S}(x_{t})).$ (4)
(2) Action Consistency Loss. For action tokens, we similarly minimize KL divergence:
$\mathcal{L}_{\text{Act}}=\sum_{t=1}^{T}m_{a}(t)\cdot\text{KL}(p_{T}(x_{t})\|p_{S}(x_{t})).$ (5)
Final Objective. The total structured loss aggregates these terms:
@@ 3.4 Structured Agent Distillation Objectives
Semantic Decoupling. While the overall loss is additive in form, our supervision is fundamentally different from flat token-level imitation. We explicitly decompose the learning signal into structurally disjoint spans—[REASON] and [ACT]—and apply segment-specific losses to each, preserving the semantics of multi-phase agent behaior.
@@
CoT-Policy Alignment Loss ($\mathcal{L}_{\text{CoT}}$) supervises predictions within the reasoning span, promoting coherent multi-step inference aligned with teacher thought patterns. The reasoning-action example below highlights span-specific KL penalties.
Action Consistency Loss ($\mathcal{L}_{\text{Act}}$) applies only to the action span, enforcing accurate replication of grounded decisions.
@@
Each token is assigned to exactly one functional span using binary masks $\{m_{r}(t),m_{a}(t)\}$, which gate gradient flow. This masking is not heuristic—it enforces semantic separation during training, ensuring that the student independently learns high-level reasoning and low-level execution.
Unlike token-level KL with soft targets, our structure-aware formulation avoids loss interference across phases, better modeling causal dependencies (e.g., reason → act). Ablation results confirm that removing either component harms performance, underscoring their complementary roles in agent compositionality.
This separation ensures that students do not overfit to output correctness alone, but also learn to emulate coherent decision logic.
Example. Reasoning-action trajectory from the teacher and the corresponding imitation by the student:
@@
Although the student executes the correct action (search[tray]), its reasoning deviates from the teacher’s thought process. The CoT-Policy Alignment Loss penalizes semantic deviations within the reasoning span ([REASON]) by computing KL divergence between the teacher and student token distributions. When the student generates thought tokens inconsistent with the teacher’s, the KL divergence increases, producing gradients that push the student toward reproducing the teacher’s multi-step reasoning trace. In contrast, the Action Consistency Loss rewards correct predictions in the [ACT] span, allowing action alignment even when reasoning differs.
3.5 Curriculum Sampling in Structured Distillation

### v3

@@
Abstract.
Large language models (LLMs) exhibit strong capabilities as decision-making agents by interleaving reasoning and actions, as seen in ReAct-style frameworks. Yet, their practical deployment is constrained by high inference costs and large model sizes. We propose Structured Agent Distillation, the first framework to distill a ReAct-based LLM agent into a smaller model while preserving both reasoning fidelity and action consistency. Our method introduces a structured, span-level distillation strategy that explicitly segments trajectories into reasoning and action spans, enabling fine-grained alignment beyond standard token-level imitation. Unlike other advanced distillation methods, Our method segments trajectories into [REASON] and [ACT] spans, applying segment-specific losses to align each component with the teacher’s behavior. This structure-aware supervision enables compact agents to better replicate the teacher’s decision process. Experiments on ALFWorld, HotPotQA-ReAct, and WebShop show that our approach consistently outperforms token-level and imitation learning baselines, achieving significant compression with minimal performance drop. Scaling and ablation results further highlight the importance of span-level alignment for efficient and deployable agents. We will release code upon acceptance.
Key words and phrases:
Agent Distillation, Span-Level Alignment, Reasoning and Action Segmentation, Large Language Models (LLMs), Span-Level Alignment
1. Introduction
@@
Despite their effectiveness, LLM-based general agents remain costly to deploy due to model size and inference overhead. To address this, recent work distills large agents into smaller student models. However, most approaches rely on token-level supervision hinton2015distilling; turc2019well; eldan2023tinystories; gu2024minillm; minixhofer2025cross; cui2025multi; zhang2024knowledge, which treats the agent trajectory as a flat token sequence and aligns predictions step by step—ignoring its structured composition of reasoning and action.
@@
Our Contributions:
  • Our work is the first to distill ReAct-style LLM agents using structured span-level supervision: we segment trajectories into reasoning and action spans and apply span-specific alignment via token-level masks, improving over naive token-level distillation.
@@
2. Motivation
2.1. Why Token-Level Distillation Falls Short
@@
Table 1. Comparison between LLMs and LLM-based General Agents in terms of token efficiency
@@
Output Format Final text sequence Task trajectory (reasoning steps + actions)
2.2. Motivation: Toward Structured Agent Distillation
@@
Table 2. Comparison of LLM agent training frameworks. Only our method supports all four dimensions of structured agent distillation. Tool: supports external API calls or tool use; R–A Align.: aligns structured reasoning and action spans; Seg.-aware Sup.: applies supervision across reasoning–action sequences; Curric.: uses trajectory difficulty for progressive training kumar2010self.
Framework Tool R–A Align. Seg.-aware Sup. Curric.
@@
3. Structured Agent Distillation Framework
Figure 1. Structured Agent Distillation framework. The teacher provides segmented reasoning–action trajectories. The student aligns CoT traces and actions via span-specific KL losses, with projected gradients and curriculum sampling for stability.
[…]
Structured Agent Distillation framework.
Structured Agent Distillation (SAD) segments teacher trajectories into reasoning (Reason) and interaction (Action/Observation) spans, each supervised independently to promote phase-specific alignment.
As shown in Figure 1, the teacher agent, given an observation and task prompt, produces [REASON]ing traces and [ACT]ion outputs, forming a trajectory $\tau=(\texttt{reason},\texttt{action})$ used for curriculum sampling.
The student learns from $\tau$ via two losses: (1) CoT-Policy Alignment $\mathcal{L}_{\text{CoT}}$ aligns reasoning, and (2) Action Consistency $\mathcal{L}_{\text{Act}}$ aligns decisions.
Refer to Appendix A for comprehensive analysis.
3.1. Problem Formulation
@@
Given a teacher policy $\pi_{T}(\tau)$, the goal is to train a compact student policy $\pi_{\theta}(\tau)$ such that
@@
$\tau^{\prime}=\texttt{[REASON]}\ r_{1}\cdots r_{k}\ \texttt{[ACT]}\ a_{1}\cdots a_{m}.$
We tokenize this as
$x=\texttt{Tokenize}(\tau^{\prime})=(x_{1},x_{2},\ldots,x_{T}),$ (3)
and assign each token $x_{t}$ a segment label $s_{t}\in\{\text{Reason},\text{Action}\}$, indicating its span. These labels are used to compute segment-aware losses during training.
For clarity, we adopt the following notation: $\tau$ denotes the structured reasoning–action trajectory, $\tau^{\prime}$ its linearized form with explicit segment markers, and $x=\texttt{Tokenize}(\tau^{\prime})$ the token sequence processed by the model. Accordingly, $\pi_{\theta}$ always operates on tokenized inputs $x$, while $\tau$ and $\tau^{\prime}$ are used only for segmentation and mask construction.
3.2. Trajectory Segmentation
@@
$(\tau^{(r)},\tau^{(a)})\leftarrow\text{Segment}(\tau),$
where $\tau^{(r)}$ denotes the reasoning span and $\tau^{(a)}$ denotes the action span. This segmentation is performed via lightweight rule-based parsing based on prompt templates consistent across tasks (see Appendix D). The segmented trajectory is then tokenized into a sequence $x$ defined in Eq. (3).
While the above formulation assumes a single reasoning–action pair, SAD naturally extends to multi-step ReAct trajectories $\tau=[(r^{(i)},a^{(i)},o^{(i)})]_{i=1}^{K}$ by applying union masks over multiple reasoning/action spans. A detailed construction and example are provided in Appendix M.
3.3. Structured Agent Distillation Objectives
To supervise student agents under SAD, we align binary token masks $m_{r}(t)$ and $m_{a}(t)$ with the tokenized sequence $x=(x_{1},\ldots,x_{T})$, enabling selective supervision over structurally distinct parts of the trajectory.
(1) CoT-Policy Alignment Loss.
For reasoning tokens, the student’s conditional distribution $p_{S}(\cdot\mid x_{<t})$ is aligned with the teacher’s distribution $p_{T}(\cdot\mid x_{<t})$ using Kullback–Leibler(KL) divergence:
$\mathcal{L}_{\text{CoT}}=\sum_{t=1}^{T}m_{r}(t)\,\mathrm{KL}\!\left(p_{T}(\cdot\mid x_{<t})\,\|\,p_{S}(\cdot\mid x_{<t})\right)$ (4)
(2) Action Consistency Loss.
For action tokens, we similarly minimize KL divergence:
$\mathcal{L}_{\text{Act}}=\sum_{t=1}^{T}m_{a}(t)\,\mathrm{KL}\!\left(p_{T}(\cdot\mid x_{<t})\,\|\,p_{S}(\cdot\mid x_{<t})\right)$ (5)
CoT-Policy Alignment operates over the full vocabulary during [REASON] spans to guide the student’s intermediate reasoning steps, encouraging alignment with the teacher’s chain-of-thought. In contrast, Action Consistency Loss applies KL over a discrete action space during [ACT] spans, ensuring that the student replicates the teacher’s action decisions.
Final Objective.
The total structured loss aggregates these terms:
@@ 3.3. Structured Agent Distillation Objectives
Clarification. Although the loss terms are unscaled, this formulation is not equivalent to computing a single token-level KL over the entire vocabulary. Token-level KL normalizes over the joint token space $\mathcal{V}_{r}\cup\mathcal{V}_{a}$, which couples gradients from frequent reasoning tokens and rare but critical action tokens. In contrast, SAD applies KL over disjoint normalization domains ($\mathcal{V}_{r}$ for reasoning, $\mathcal{V}_{a}$ for action), thereby changing both the normalization space and gradient direction. This decomposition alters the optimization geometry and prevents cross-span interference, constituting a fundamental difference from a flat token-level KL even when $\lambda_{r}$ and $\lambda_{a}$ are equal. While our formulation assumes teacher-forced alignment during training, the loss can be extended to accommodate mismatched trajectories via alignment-based matching, as detailed in Appendix H.
3.4. Optimization View: Gradient Projection and Span Decoupling
Figure 2. Optimization structure of Structured Agent Distillation (SAD) via gradient projection. Standard token-level KL (gray) couples reasoning and action gradients, leading to a conflict angle $\theta$. SAD resolves this by projecting gradients onto reasoning ($\mathcal{V}_{r}$) and action ($\mathcal{V}_{a}$) subspaces, forming an orthogonal decomposition. The resulting projected update (green) follows the composite direction $\Pi_{\text{reason}}(\nabla_{\theta}\mathcal{L}_{\text{CoT}})+\Pi_{\text{action}}(\nabla_{\theta}\mathcal{L}_{\mathrm{Act}})$, eliminating cross-span interference and clarifying SAD’s optimization geometry.
[…]
Diagram showing geometric interpretation of Structured Agent Distillation (SAD), illustrating gradient projections and subspaces.
To better understand the functional benefit of SAD, we present an optimization-based interpretation of how span-specific KL losses reshape the gradient landscape. Rather than assuming a cognitive or semantic split, SAD introduces an optimization structure that avoids interference between reasoning and action supervision signals.
As illustrated in Figure 2, a token-level KL couples heterogeneous gradients into a single update direction (gray arrow), creating a conflict angle $\theta$ between reasoning and action forces in parameter space. SAD reformulates this process as an orthogonal gradient projection: reasoning and action components are separately normalized in their respective subspaces ($\mathcal{V}_{r}$, $\mathcal{V}_{a}$) and then recombined geometrically. This projection removes cross-span interference and yields span-specific updates, altering the overall optimization geometry.
Let $\mathcal{V}_{r}$ and $\mathcal{V}_{a}$ denote the token domains for reasoning and action, along with $\mathcal{V}_{r}\cap\mathcal{V}_{a}=\varnothing$. A standard token-level KL minimizes
$\nabla_{\theta}\mathcal{L}_{\text{token}}=\nabla_{\theta}\,\mathrm{KL}\!\left(p_{T}(\cdot\mid x_{<t})\,\|\,p_{S}(\cdot\mid x_{<t})\right),$ (7)
where normalization over the entire vocabulary $\mathcal{V}_{r}\cup\mathcal{V}_{a}$ forces shared probability mass between semantically incompatible tokens. This coupling biases gradients toward frequent reasoning tokens and suppresses rare but task-critical action tokens, explaining why a single distribution cannot disentangle these behaviors.
SAD resolves this by restricting the KL computation to disjoint subspaces:
$\nabla_{\theta}\mathcal{L}_{\text{SAD}}=m_{r}(t)\,\nabla_{\theta}\mathrm{KL}_{\mathcal{V}_{r}}\!\left(p_{T}\,\|\,p_{S}\right)+m_{a}(t)\,\nabla_{\theta}\mathrm{KL}_{\mathcal{V}_{a}}\!\left(p_{T}\,\|\,p_{S}\right),$ (8)
which performs gradient projection onto reasoning and action subspaces,
$\nabla_{\theta}\mathcal{L}_{\text{SAD}}=\Pi_{\text{reason}}\!\big(\nabla_{\theta}\mathcal{L}_{\text{CoT}}\big)+\Pi_{\text{action}}\!\big(\nabla_{\theta}\mathcal{L}_{\mathrm{Act}}\big).$ (9)
This projection changes both the normalization domain and gradient direction, eliminating cross-span interference and yielding span-specific updates. Hence, the distinction from token-level KL is geometric, not cosmetic: SAD introduces a structure-aware gradient decomposition rather than simply applying the KL divergence to a smaller subset of tokens.
3.5. Multi-Step ReAct Trajectories and Multi-Span Masks
Trajectory Model. Extending the single-step formulation in Eq. (1), we generalize SAD to multi-step ReAct episodes composed of alternating reasoning, action, and observation segments:
$\tau=\big[(r^{(1)},a^{(1)},o^{(1)}),(r^{(2)},a^{(2)},o^{(2)}),\ldots,(r^{(K)},a^{(K)},o^{(K)})\big],$ (10)
where $r^{(i)}$, $a^{(i)}$, and $o^{(i)}$ denote the reasoning trace, executed action, and subsequent observation at step $i$. Linearization yields
$\tau^{\prime}=\prod_{i=1}^{K}\big(\texttt{[REASON]}\penalty 10000\ r^{(i)}\penalty 10000\ \texttt{[ACT]}\penalty 10000\ a^{(i)}\penalty 10000\ \texttt{[OBS]}\penalty 10000\ o^{(i)}\big),\qquad x=\texttt{Tokenize}(\tau^{\prime}).$
This formulation extends SAD beyond the single-step assumption, enabling structured supervision across multiple reasoning–action cycles.
Segment-Aware Mask Construction. Each token $x_{t}$ in the sequence is assigned a binary membership mask:
$m_{r}(t)$ $=\mathbf{1}[x_{t}\in\cup_{i}r^{(i)}],m_{a}(t)$ $=\mathbf{1}[x_{t}\in\cup_{i}a^{(i)}],$ (11)
$m_{o}(t)$ $=\mathbf{1}[x_{t}\in\cup_{i}o^{(i)}],$
We enforce an exactly-one constraint for every token:
$m_{r}(t)+m_{a}(t)+m_{o}(t)=1,\quad\forall t\in[1,T],$ (12)
ensuring non-overlapping and functionally disjoint spans. Reasoning and action masks ($m_{r},m_{a}$) are used for supervision, while observation masks ($m_{o}$) indicate environmental feedback.
Observation Handling. Each token is assigned to exactly one of the three functional categories—reasoning, action, or observation—ensuring disjoint span boundaries. However, only reasoning and action tokens contribute to the distillation loss. Observation tokens ($m_{o}(t)=1$) are excluded from the distillation loss, as they encode deterministic feedback from the environment rather than agent behavior. This exclusion prevents the student from overfitting to static textual observations and focuses learning on reasoning and decision quality. Nonetheless, the framework permits optional extensions: (1) adding an auxiliary cross-entropy term for perceptual grounding, or (2) defining a separate observation head for multimodal tasks. All reported experiments adopt the exclusion setting.
Supervision over Multi-Span Masks. Structured losses defined in Section 3.3 (Eq. (4) and (5)) are applied independently to each reasoning and action span, as detailed in Appendix L. with $m_{r}$ and $m_{a}$ computed as the union of all [REASON] and [ACT] segments. This union-mask construction generalizes SAD to multi-turn reasoning–acting trajectories, preserving disjoint functional roles and preventing cross-span gradient interference.
Illustrative Example. A two-step episode (K=2K{=}2) from ALFWorld: [REASON] I will first check the table. [ACT] search[tray] [OBS] You see a tray.
[REASON] Now I will pick it up. [ACT] pickup[tray] [OBS] Tray in inventory.
Here, reasoning tokens correspond to $m_{r}{=}1$, action tokens to $m_{a}{=}1$, and observation tokens to $m_{o}{=}1$. Only reasoning and action tokens receive gradient updates, clarifying span-level supervision semantics and aligning with multi-turn ReAct agent behavior.
3.6. Optimization Analysis and Intuition
Why it works in practice. The decoupled objectives alter both gradient magnitude and direction across spans. Empirically, SAD reduces step-to-step gradient variance and improves training stability under limited data. This acts as an implicit curriculum: reasoning spans provide dense signals for high-level coherence, while action spans yield sparse but decisive grounding signals. The balanced supervision accelerates convergence and strengthens reasoning–action compositionality.
SAD provides a principled change in optimization geometry that (1) moves beyond simplified cognitive assumptions, (2) explains why a single token-level distribution fails to separate reasoning and action, and (3) establishes a structure-aware, theoretically grounded distillation objective.
3.7. Semantic Decoupling and Example
@@
While the overall loss is additive in form, our supervision is fundamentally different from flat token-level imitation. We explicitly decompose the learning signal into structurally disjoint spans—[REASON] and [ACT]—and apply segment-specific losses to each, preserving the semantics of multi-phase agent behavior.
@@
CoT-Policy Alignment Loss ($\mathcal{L}_{\text{CoT}}$) supervises predictions within the reasoning span, promoting coherent multi-step inference aligned with the teacher’s thought patterns. Action Consistency Loss ($\mathcal{L}_{\text{Act}}$) applies only to the action span, enforcing the accurate replication of grounded decisions.
@@
Each token is assigned to exactly one functional span using binary masks $\{m_{r}(t),m_{a}(t)\}$, which gate gradient flow. This masking enforces semantic separation during training, ensuring that the student independently learns high-level reasoning and low-level execution. Unlike token-level KL with soft targets, our structure-aware formulation avoids loss interference across phases, better modeling causal dependencies (reason $\rightarrow$ act). Ablations confirm that removing either component harms performance.
Example.
@@
Although the student executes the correct action, their reasoning deviates from the teacher’s thought process. $\mathcal{L}_{\text{CoT}}$ penalizes semantic deviations within [REASON] via KL divergence between teacher and student token distributions, producing gradients that align multi-step reasoning. $\mathcal{L}_{\text{Act}}$ rewards correct predictions in [ACT], allowing action alignment even when reasoning differs.
3.8. Curriculum Sampling in Structured Distillation